OpenStudy (anonymous):

A rock falls from a cliff 125 meters above the ground. The equation h=125-5t(^2) represents the height (h) of the rock after (t) seconds. How many seconds does it take for the rock to land on the ground?

OpenStudy (anonymous):

is it approximately 0.45??

OpenStudy (anonymous):

I forgot my Algebra 2 years ago after I completed the class. :p Maybe someone else can run the problem and check the answer?

OpenStudy (anonymous):

lol...ohkeyy!

OpenStudy (anonymous):

We are putting h=0, right?

OpenStudy (anonymous):

no...bt it will get o...as 125-125

OpenStudy (anonymous):

h=0, as the distance is measure from the ground, so once it reaches the ground h=0 thus 0 = 125 -5t^2 t^2=25 t=5

OpenStudy (anonymous):

plus/minus 5 but the minus figure doesn't have any meaning here...

OpenStudy (anonymous):

Clarifying; it will take 5 seconds for it to come to the ground? Or the original .45

OpenStudy (anonymous):

The equation h=125-5t(^2) represents the height (h) of the rock after (t) seconds.

OpenStudy (anonymous):

i guess majority is winning here...though m nt convinced

OpenStudy (anonymous):

Ok, so 5. Thanks guy :3